The density of copper is 8.9*10^3 kg/m^3, and each copper atom has a mass if 63u (1u=1.6605*10^-27kg.) Estimate the average distance between the centers of the neighboring copper atom?
1
Expert's answer
2019-10-15T06:14:19-0400

l=(ρ*1000*Na/M)-1/3


I=(8.9*103*1000/63)*6.022*1023)-1/3=22.7*10-9 m = 22.7 nm.


I - distance between atoms

ρ - density

M - molar mass

Na – Avogadro number